OpenStudy (anonymous):

calculate the wavelength of emr that has a frequency of 3.0*10^5 hz

OpenStudy (anonymous):

c= speed of light = 3 x10^8 m/s c = f L = (frequency ) (wavelength) L = c/F =(3x10^8 m/s) / (3 x 10^5 /s) = 1000m